### v1

# Images are Achilles’ Heel of Alignment: Exploiting Visual Vulnerabilities for Jailbreaking Multimodal Large Language Models

Abstract
In this paper, we study the harmlessness alignment problem of multimodal large language models (MLLMs). We conduct a systematic empirical analysis of the harmlessness performance of representative MLLMs and reveal that the image input poses the alignment vulnerability of MLLMs. Inspired by this, we propose a novel jailbreak method named HADES, which hides and amplifies the harmfulness of the malicious intent within the text input, using meticulously crafted images. Experimental results show that HADES can effectively jailbreak existing MLLMs, which achieves an average Attack Success Rate (ASR) of 90.26% for LLaVA-1.5 and 71.60% for Gemini Pro Vision. Our code and data will be publicly released.
Warning: this paper contains example data that may be offensive.
Keywords:
Multimodal Large Language Models Harmlessness Alignment Adversarial Attack
1 Introduction
[…]
Recently, by leveraging the powerful capacity of large language models (LLMs) [37], a variety of multimodal large language models (MLLMs) [36] have ${}^{{\ddagger}}$, which can process both textual and visual information similarly as that LLMs process textual input. MLLMs have not only shown superior performance in various visual-language tasks but also possess the capability to engage in image-related dialogues with human users [18, 39]. However, MLLMs also confront similar harmlessness challenges that afflict their backbone LLMs.
Despite undergoing harmlessness alignment like reinforcement learning from human feedback (RLHF) [23], LLMs remain vulnerable to black-box attacks (e.g., sophisticated jailbreak prompts [5]) or white-box attacks (e.g., gradient-based adversarial inputs [40]). Since MLLMs are generally built on top of existing LLMs, they inevitably suffer from these safety issues. To study the harmlessness alignment of MLLMs, recent work either evaluates the harmlessness of MLLMs in response to harmful instructions [19, 31], or assesses the model robustness by utilizing adversarial images [24, 12]. These studies suggest that the integration of the visual modality might exacerbate safety concerns for MLLMs compared to their backbone LLMs. As illustrated in Fig. 1, even one of the state-of-the-art MLLMs, Gemini Pro Vision [1], would be influenced by the blank or harmful images, leading to improper outputs with harmful words.
Figure 1: An example to show the influence of the visual modality on harmlessness alignment of Gemini Pro Vision. The harmful information is highlighted in red.
However, it still lacks of deep understanding of how safety concerns occur in MLLMs and how they might be different from those in LLMs. Considering this issue, this work aims to systematically analyze the sourcing factors that violate the harmlessness alignment of MLLMs. We conduct detailed empirical studies on representative MLLMs, specifically investigating their performance on harmful instructions accompanied by images. Our findings are threefold: (1) Images can be backdoors for the harmlessness alignment of MLLMs. The inclusion of images in the input can significantly increase the harmfulness ratio of MLLMs’ outputs; (2) Cross-modal finetuning undermines the alignment abilities of the backbone LLM for a given MLLM. The more parameters that are finetuned, the more severe the disruption is; (3) The harmfulness of MLLMs’ responses is positively correlated with the harmfulness of the image content. These findings reveal that the visual modality introduces additional alignment vulnerabilities in MLLMs, which can be exploited to further jailbreak these models.
Motivated by these empirical findings, we propose a novel jailbreak approach called HADES, standing for Hiding and Amplifying harmfulness in images to DEStroy multimodal alignment, to assess the adversarial robustness of both open- and closed-source MLLMs. Specifically, our approach introduces a three-stage attack strategy. First, it extracts the harmful information from the text input into typography and replaces such text with a text-to-image pointer, which guides the model to focus on image information. In this way, we transfer the harmful input from the well-aligned text side into the image side, inducing models to be more prone to generate harmful outputs. Second, HADES attaches another harmful image to the original typography. This image is created by an image generation model, and its harmfulness has been amplified for multiple turns via prompt optimization. Third, HADES optimizes an adversarial noise via gradient update, towards inducing the MLLM to follow harmful instructions. The learned noise will be integrated into the previous image for jailbreaking MLLMs.
In summary, our main contributions are as follows:
  • $\bullet$
    We conduct detailed empirical studies on the harmfulness alignment of representative MLLMs, and systematically investigate the possible sourcing factors that violate the harmfulness alignment of MLLMs. The results reveal that the visual modality of MLLMs poses a critical alignment vulnerability.
  • $\bullet$
    We introduce HADES, a novel jailbreak approach that hides and amplifies the harmfulness of the original malicious intent using meticulously crafted images. Experimental results show that both open-source MLLMs based on aligned LLMs and powerful closed-source MLLMs struggle to resist HADES. Notably, HADES achieves an Attack Success Rate (ASR) of 90.26% on LLaVA-1.5 and 71.60% on Gemini Pro Vision.
2 Empirical Harmlessness Analyses of MLLMs
In this section, we conduct a systematic investigation to examine whether and how visual input influences the harmlessness alignment of MLLMs. We first introduce the data collection process in Sec. 2.1 and the evaluation settings in Sec. 2.2, then evaluate mainstream open- and closed-source MLLMs in Sec. 2.3.
2.1 Evaluation Data Collection
To evaluate the harmlessness alignment of MLLMs, we collect a dataset comprising 750 harmful instructions across 5 scenarios. Each instruction includes a harmful keyword or key phrase and is paired with a harmful image related to the keyword or key phrase. We present the collection process below and show the pipeline in appendix.
First, based on existing harmful scenarios of LLMs [14], we select five representative ones that are related to the visual information in the real world: (1) Violence, Aiding and Abetting, Incitement; (2) Financial Crime, Property Crime, Theft; (3) Privacy Violation; (4) Self-Harm; and (5) Animal Abuse. For simplicity, these categories are referred to as Violence, Financial, Privacy, Self-Harm, and Animal, respectively. Next, we adopt GPT-4 to generate 50 keywords for each of the above harmful categories, and then synthesize three harmful but distinct instructions based on each keyword. The prompts employed for the above process are documented in appendix. In this way, we guarantee that each instruction includes only one harmful element (i.e., the keyword or the key phrase), which could be accurately depicted by an image. Thus, we can pair each instruction with a corresponding real-world image that is relevant to the harmful keyword/phrase. Specifically, we first retrieve five images from Google using each keyword/phrase as the query, and then employ CLIP ViT-L/14 [25] to select the image that best matches the semantic representation of the keyword/phrase.
2.2 Evaluation Settings
We evaluate both representative open-source (i.e., LLaVA-1.5 [17], LLaVA-1.5 LoRA (represented by LLaVA-1.5L), MiniGPT-v2 [6] and MiniGPT-4 [39]) and powerful closed-source MLLMs (i.e., Gemini Pro Vision [1], short as Gemini ProV) and GPT-4V [22]) on our dataset. To discuss the influence of the visual modality on harmlessness alignment, we design four evaluation settings:
  • $\bullet$
    Backbone: We evaluate the backbone LLMs of MLLMs on harmful instructions, which have not been fine-tuned with cross-modal data, i.e., Vicuna v1.5-7B [8] for LLaVA-1.5, LLaMA-2-Chat-7B [30] for MiniGPT-4 and MiniGPT-v2, GPT-4 for GPT-4V, and Gemini Pro for Gemin ProV.
  • $\bullet$
    Text-only: We evaluate MLLMs purely on harmful instructions.
  • $\bullet$
    Blank: We evaluate MLLMs on harmful instructions paired with a 500 $\times$ 500 blank image.
  • $\bullet$
    Toxic: We evaluate MLLMs on harmful instructions paired with the previously selected harmful images (described in Sec. 2.1).
We employ Attack Success Rate (ASR) as the evaluation metric of harmlessness alignment, which is calculated by:
$\text{ASR}=\frac{\sum_{i=1}^{N}\mathbbm{1}_{\{\mathcal{J}(y_{i})=\text{True}\}}}{N},$ (1)
where $y_{i}$ is the model’s response, $\mathbbm{1}$ is an indicator function that equals to 1 if $\mathcal{J}(y_{i})=\text{True}$ and 0 otherwise, $N$ is the total number of instructions and $\mathcal{J}$ is the harmfulness judging model, outputting True or False to indicate whether $y_{i}$ is harmful. We adopt Beaver-dam-7B [14] as $\mathcal{J}$, which has been trained on high-quality human feedback data about the above harmful categories.
Table 1: The evaluation results of representative MLLMs on the dataset we collected. (Train) represents the cross-modal finetuning strategies of MLLMs. Animal, Financial, Privacy, Self-Harm, and Violence represent the ASR of MLLMs on instructions from these categories. Average represents the average ASR across all categories. $+$ and -- represents the change of ASR compared to the Backbone setting.
Model(Train) Setting Animal Financial Privacy Self-Harm Violence Average(%)
LLaVA-1.5(Full) Backbone 17.33 46.00 34.67 12.00 34.67 28.93
Text-only 22.00 40.00 28.00 10.00 30.67 26.13($-\phantom{0}2.80$)
Blank 38.00 66.67 68.00 30.67 67.33 54.13($+25.20$)
Toxic 54.00 77.33 82.67 46.67 80.00 68.13($+39.20$)
LLaVA-1.5L(LoRA) Backbone 17.33 46.00 34.67 12.00 34.67 28.93
Text-only 23.33 40.00 30.00 9.33 30.67 26.67($-\phantom{0}2.26$)
Blank 41.33 67.33 63.33 25.33 61.33 51.73($+22.80$)
Toxic 48.67 71.33 74.67 43.33 76.00 62.80($+33.87$)
MiniGPT-v2(LoRA) Backbone 0.00 0.00 0.00 0.00 0.67 0.13
Text-only 7.33 12.00 8.67 0.00 15.33 08.67($+\phantom{0}8.54$)
Blank 26.00 46.67 40.00 16.00 41.33 34.00($+33.87$)
Toxic 37.33 60.67 50.00 27.33 44.00 43.87($+43.74$)
MiniGPT-4(Frozen) Backbone 0.00 0.00 0.00 0.00 0.67 00.13
Text-only 5.33 2.67 1.33 1.33 3.33 02.80($+\phantom{0}2.67$)
Blank 15.33 13.33 6.67 0.00 8.67 08.80($+\phantom{0}8.67$)
Toxic 28.67 35.33 18.67 9.33 25.33 23.47($+23.34$)
Gemini ProV(-) Backbone 1.70 13.80 12.08 1.20 8.70 7.50
Text-only 0.00 0.00 0.00 0.00 0.00 0.00($-\phantom{0}7.50$)
Blank 13.33 42.67 34.00 5.33 21.33 23.33($+15.83$)
Toxic 19.33 52.00 45.33 6.67 30.00 30.67($+23.17$)
GPT-4V(-) Backbone 0.00 2.00 2.67 0.00 0.67 1.07
Text-only 1.33 8.67 6.00 0.67 7.33 4.80($+\phantom{0}3.73$)
Blank 2.00 4.67 6.00 0.00 6.67 3.87($+\phantom{0}2.80$)
Toxic 2.00 14.00 14.00 0.00 6.00 7.20($+\phantom{0}6.13$)
2.3 Evaluation Results
The evaluation results are presented in Tab. 1. We list the ASR results of 5 harmful scenarios under 4 evaluation settings and calculate the average ASR across all scenarios. From the results, we can summarize three major findings:
Images can be alignment backdoors of MLLMs. When comparing the performance of each model under the Backbone and Text-only settings, the harmlessness alignment of MLLMs does not significantly deviate from that of their backbone LLMs, and even exhibit enhanced defense capability, e.g., LLaVA-1.5 (--2.80%) and Gemini ProV (--9.65%). However, once adding images, regardless of whether their contents are harmful or not, the ASR results of MLLMs would be greatly improved even under the Blank setting using harmless images, e.g., LLaVA-1.5 ($+$25.20%) and MiniGPT-v2 ($+$33.87%). It indicates that images can be the alignment backdoor of MLLMs, which would undermine MLLMs’ capability of defending against harmful text input.
More parameters tuned, less alignment left. By examining the performance of open-sourced MLLMs, we notice that their alignment performance is highly relevant to their training strategies during cross-modal finetuning. Generally, for the same backbone model, the more parameters are optimized during finetuning, the harmlessness alignment would be affected more significantly. For example, under the Toxic setting, MiniGPT-4, with the fixed backbone LLM, is more robust than the LoRA-finetuned MiniGPT-v2, achieving lower ASR results (23.47% vs. 43.87%). Similarly, the full-parameter finetuned LLaVA-1.5 also generates more harmful responses than LoRA-finetuned LLaVA-1.5L (68.13% vs. 62.80%). The reason may be that the cross-modal finetuning process would hurt the harmlessness alignment of the backbone LLMs.
Harmful images are more likely to elicit harmful outputs. We observe that MLLMs are more prone to produce harmful outputs when presented with harmful images. It holds for both open-source and closed-source models, as their ASR results under the Toxic setting greatly exceed the results from all other settings, e.g., MiniGPT-4 (23.47% vs. 8.80%) and LLaVA-1.5 (68.13% vs. 54.13%). It indicates that it is hard for current MLLMs to defend against harmful image inputs. With the increase in image harmfulness, MLLMs might be increasingly prone to generate more harmful outputs.
Figure 2: Given a harmful textual instruction, HADES involves a three-step procedure: (1) removes the harmful content from the text into typography; (2) combines it with a harmful image generated by a diffusion model, using an iteratively refined prompt from an LLM; (3) appends an adversarial image on top of the image, which elicits the MLLM to generate affirmative responses for harmful instructions.
3 The Proposed Jailbreak Approach: HADES
Based on empirical studies, the visual input potentially brings vulnerabilities to the harmlessness alignment of MLLMs. However, it is not easy to manually craft massive adversarial samples that can successfully jailbreak MLLMs, which help detect the weak spots in real-world applications. To address it, based on harmful textual instructions, we propose a novel method to automatically synthesize high-quality adversarial examples, called HADES that stands for Hiding and Amplifying harmfulness in images to DEStroy multimodal alignment.
Typically, an MLLM is composed of an LLM $\mathcal{M}$, an image encoder E and a projection layer W. The generation process of MLLMs can be formulated as:
$y=\mathcal{M}([\textbf{W}\cdot\textbf{E}(i),t]),$ (2)
where $i$ and $t$ are the input image and text, and $y$ is the model’s output. Given a harmful $t$, HADES aims to modify $t$ to $t^{\prime}$ by adding a text-to-image pointer and crafts harmful images $i^{\prime}$. Thereby, HADES transfers the malicious intent to the less-aligned image side of the MLLM, inducing it to generate harmful responses $y^{\prime}$. The whole process of HADES is presented in Fig. 2.
3.1 Hiding Harmfulness from Text to Image
By training on human preference data, existing LLMs learn to align with human values and refuse to respond to harmful text inputs. MLLMs derived from these LLMs naturally inherit the defense capacity for text inputs but leave the image side vulnerable to harmful content. Hence, we propose transferring harmful information from the well-aligned text side to the less-aligned image side, to bypass the defense mechanisms of MLLMs.
Specifically, we replace the harmful keyword or key phrase from each text instruction $t$ with a text-to-image pointer and utilize images to represent it. On the text side, we categorize all the keywords into three classes: objects, concepts, or behaviors. For the keywords falling under the first two categories, the text-to-image pointer is “the object/concept in the image”, while for the keywords denoting behaviors, the text-to-image pointer is “conduct the behavior on”. On the image side, as the keywords may represent abstract concepts or behaviors that are difficult for models to grasp when depicted by real-world images, we employ typography to represent these keywords. As a result, the generation process of MLLMs can be formally given as:
$y=\mathcal{M}([\textbf{W}\cdot\textbf{E}(i_{\text{typ}}),t^{\prime}]),$ (3)
where $i_{\text{typ}}$ is the typography of the keyword and $t^{\prime}$ is the modified instruction. By doing so, $t^{\prime}$ no longer explicitly contains any harmful information, yet models can still infer the original harmful intent by referring to $i_{\text{typ}}$.
3.2 Amplifying Image Harmfulness with LLMs
Our empirical study reveals that when the image input becomes more harmful, MLLMs also tend to generate more harmful responses. Therefore, we propose to append a harmful image to the previous typography $i_{typ}$ to amplify their harmfulness. Since the harmfulness of real-world images is always limited, we introduce diffusion models as the harmful image generator. In addition, we utilize LLMs as the attacker model to iteratively optimize the prompt for diffusion models to further increase the harmfulness of generated images.
The whole procedure of image harmfulness optimization is presented in appendix. We leverage the harmfulness of the caption as the proxy for the image’s harmfulness, as the harmfulness of text is easier to quantify than images. We consider an iterative process to generate harmful images. We first ask ChatGPT to modify the original instruction $t$ into an initial image generation prompt $p_{0}$ and generate an initial image. At step $k$, the caption model $\mathcal{C}$ generates a caption $c_{k}$ for the target image $i_{\text{opt}}^{k}$. Subsequently, the judging model $\mathcal{J}$ assesses the image’s harmfulness with a score, $s_{k}$, on a scale from 1 to 10, where a higher score indicates greater harmfulness. $\mathcal{J}$ also explains the reason for its score in $exp_{k}$. All this information ($p_{k}$, $c_{k}$, $s_{k}$, and $exp_{k}$) is appended to the conversation history $h$ and sent to the attacker model $\mathcal{A}$. which first suggests improvements to the prompt and then generates the refined image generation prompt $p_{k+1}$. The refined prompt is then used to generate a new image $i_{\text{opt}}^{k+1}$. The above process circulates until reaching the maximum iteration turn $K$, which is set to 5. In practice, we choose GPT-4-0613 as both the attacker and judging model by utilizing different system prompts (as presented in the appendix). For captioning and image generation, we adopt LLaVA-1.5 and PixArt-$\alpha$ [7], respectively.
The optimized images $i_{\text{opt}}$ is then vertically concatenated with the previous typography $i_{\text{typ}}$, which can be formulated as:
$y=\mathcal{M}([\textbf{W}\cdot\textbf{E}(i_{\text{opt}}\oplus i_{\text{typ}}),t^{\prime}]).$ (4)
In this way, both images can mutually enhance their respective attack effects. The utilization of $t^{\prime}$ forces the model to focus more on the image, thereby increasing its sensitivity to harmful content. Simultaneously, $i_{\text{opt}}$, which is semantically close to the original instructions, serves as the additional context and aids the model in understanding the original harmful intent of masked instructions, which can partially compensate for some models’ insufficient OCR capabilities.
3.3 Amplifying Image Harmfulness with Gradient Update
Existing work has proven the effectiveness of adversarial images in jailbreaking MLLMs. To further enhance the attack effectiveness, HADES incorporates adversarial images by concatenating them with the previous images.
The image input of MLLMs is encoded through an image encoder and then mapped into the LLM representation space via a projection layer. Consequently, images are represented as a sequence of distinct visual tokens, collectively forming a visual prompt. It allows us to view adversarial images, those that prompt MLLMs to produce malicious responses, as text jailbreak prompts such as “Do Anything Now” [28]. Jailbreak prompts are commonly placed at the beginning of malicious instructions without compromising the models’ understanding of these instructions. Given the similarity between adversarial images and jailbreak prompts, we suggest that concatenating adversarial images on top of the image input of MLLMs can also prompt them to produce harmful information without compromising their understanding of the original image content. Consequently, the full formulation of HADES can be given as follows:
$y=\mathcal{M}([\textbf{W}\cdot\textbf{E}(i_{\text{adv}}\oplus i_{\text{opt}}\oplus i_{\text{typ}}),t^{\prime}]).$ (5)
To ensure the generalization capability of our method, we generate a single adversarial image $i_{\text{adv}}$ for all harmful instructions within the same category in our dataset. Initially, $i_{\text{adv}}$ is set to be a blank image $i_{\text{initial}}$. Subsequently, we collect a set of affirmative responses, $Y_{a}$, containing 10 examples (e.g., “Sure!” or “I can answer the question for you.”) intended as the desired output of MLLMs. We provide the target model with one harmful instruction $t_{j}$ and the adversarial image $i_{\text{adv}}$, and then select one affirmative response as the target label and compute the cross-entropy loss between the model’s output and this target. The gradient derived from this loss is then utilized to iteratively refine the adversarial image $i_{\text{adv}}$. The optimization procedure can be formulated as below:
$i_{\text{adv}}\longleftarrow i_{\text{initial}}+\operatorname*{arg\,min}_{\delta}\sum_{j=1}^{m}-\log\bigg{(}p_{\theta}(y_{j}|t_{j},i_{\text{initial}}+\delta)\bigg{)},$ (6)
where $y_{j}\in Y_{a}$, and $p_{\theta}$ represents the conditional probability generated by the target MLLM. Additionally, to maintain $i_{\text{adv}}$ as a valid image, we constrain $i_{\text{initial}}+\delta\in\mathcal{B}$ during the optimization, where $\mathcal{B}=[0,1]^{w\times h\times c}$ and $w$, $h$, and $c$ denote the width, height, and channels of $i_{\text{adv}}$, respectively.
4 Experiment
4.1 Experimental Setup
For closed-source models, we select GPT-4V and Gemini ProV. For open-source models, we also select the MLLMs used before, i.e. LLaVA-1.5 and LLaVA-1.5L. We also select LLaVA built on Llama-2-Chat-7b, as it has experienced safety RLHF. Given the limited OCR capabilities of the open-source MLLMs, they could misinterpret the keywords in $i_{\text{typ}}$. Thus, we continually prompt open-source MLLMs until they either explicitly generate the harmful keywords or reach the maximum allowed retries (set to 5 in practice). To verify the effectiveness of each component of HADES, we design four evaluation settings:
  • $\bullet$
    Typ image: Evaluate all models with the original instructions $t$ and corresponding typography $i_{\text{typ}}$.
  • $\bullet$
    $+$Text-to-image pointer: Evaluate all models with modified instructions $t^{\prime}$ and the typography $i_{\text{typ}}$. The generation process is the same as Eq. 3.
  • $\bullet$
    $+$Opt image: Evaluate all models with $t^{\prime}$ and the concatenation of $i_{\text{typ}}$ and $i_{\text{opt}}$. The generation process is the same as Eq. 4.
  • $\bullet$
    $+$Adv image: The full version of HADES. Since we don’t have access to the parameters of Gemini ProV and GPT-4V, we only evaluate open-source models with $t^{\prime}$ and the concatenation of $i_{\text{typ}}$, $i_{\text{opt}}$ and $i_{\text{adv}}$. The generation process is the same as Eq. 5.
Table 2: The evaluation results of MLLMs on instructions and images processed by HADES. T2I pointer represents Text-to-image pointer. $+$ and -- represents the change of ASR compared to the Typ image setting.
Model Setting Animal Financial Privacy Self-Harm Violence Average(%)
LLaVA-1.5 Typ image 48.67 81.33 78.00 38.67 81.33 65.60
[…]
$+$Opt image 67.33 84.00 85.33 62.00 94.00 78.53($+$12.93)
$+$Adv image 83.33 89.33 94.67 89.33 94.67 90.26($+$24.66)
LLaVA-1.5L Typ image 50.00 71.33 74.67 35.33 79.33 62.13
$+$T2I pointer 30.67 53.33 59.33 24.67 72.00 48.00(--14.13)
$+$Opt image 72.00 82.67 86.67 61.33 92.00 78.93($+$16.80)
$+$Adv image 83.33 91.33 92.67 84.67 92.67 88.93($+$26.80)
LLaVA Typ image 20.67 53.33 33.33 8.00 40.00 31.07
[…]
$+$Opt image 51.33 74.00 78.00 41.33 80.00 64.93($+$33.86)
$+$Adv image 76.00 89.33 84.67 75.33 87.33 82.53($+$51.46)
Gemini ProV Typ image 30.00 56.00 46.67 17.33 22.00 34.40
$+$T2I pointer 65.33 64.00 58.00 34.67 34.67 51.33($+$16.93)
$+$Opt image 67.33 86.67 81.33 44.00 78.67 71.60($+$37.20)
GPT-4V Typ image 0.67 1.33 4.00 0.00 2.67 1.73
[…]
$+$Opt image 2.67 24.67 27.33 1.33 19.33 15.07($+$13.34)
4.2 Experiment Results
The evaluation results in Tab. 2 demonstrate that HADES significantly enhances the attack success rate (ASR) for both open- and closed-source MLLMs. Specifically, the average ASR for the three models in the LLaVA series exceeds 80 percent. Gemini ProV also struggles to counteract harmful instructions generated by HADES with an average ASR of 71.60%. Among the evaluated models, GPT-4V exhibits the strongest defense capacity against HADES, yielding a 15.07% proportion of harmful responses. When examining the models’ performance across different categories of harmful instructions, it can be observed that they generally exhibit stronger defenses against instructions related to Animal and Self-Harm, while instructions about Financial, Privacy, and Violence categories are more likely to break through the models’ safeguards.
Under the $+$Text-to-image pointer setting, we observe diverse attack outcomes across different models. The ASR increases for LLaVA ($+$6.53%), Gemini ProV ($+$16.93%) and GPT-4V ($+$1.47%), while decreases for LLaVA-1.5 (--7.47%) and LLaVA-1.5L (--14.13%). We attribute the ASR drop on these models to two main reasons. Firstly, some models have limited OCR capabilities, which causes them to misunderstand certain instructions as benign and consequently generate harmless responses. Notably, Gemini ProV and GPT-4V, which possess advanced OCR capabilities, exhibit more harmful behavior under this setting. This suggests as the development of MLLMs, their enhancing OCR capabilities will correspondingly increase the effectiveness of our method. Secondly, the text-to-image pointer is designed to bypass the defense mechanisms of MLLMs on the text side. Therefore, its effectiveness is constrained on models with inadequate harmlessness alignment (e.g., Vicuna v1.5 in LLaVA-1.5). In such cases, the improvement in ASR could be offset by the decrease due to misunderstanding.
The incorporation of $i_{\text{opt}}$ under the $+$Opt image setting significantly increases the ASR of all models, even more than 30 percent (e.g., LLaVA ($+$33.86%) and Gemini ProV ($+$37.20%)). These results further verify our previous empirical finding: harmful images tend to elicit harmful responses. Moreover, $i_{\text{opt}}$ crafted in Sec. 3.2 helps mitigate the misunderstanding issues observed with LLaVA-1.5 and LLaVA-1.5L. The ASR of these two models increases notably compared to the $+$Text-to-image pointer setting. We attribute such results to that $i_{\text{opt}}$ always describes scenarios relevant to the instruction contents, which provides extra context that helps MLLMs to accurately understand the instruction.
Finally, by combining $i_{\text{adv}}$ under the $+$Adv image setting, the full version of HADES further increases the ASR on open-source models. Even LLaVA, whose backbone LLM is well-aligned by RLHF, achieves an average ASR of 82.53%. HADES also demonstrates promising ASR on categories that are relatively harder to jailbreak under previous settings. For LLaVA, the ASR on the Animal category rises from 51.33% to 76.00%, while the ASR on the Self-Harm category rises from 41.33% to 75.33%.
4.3 Further Analyses
In this part, we further discuss the effectiveness of our proposed approach, from image harmfulness optimization, attack transferability, and jailbreak cases.
4.3.1 Effectiveness of Image Harmfulness Optimization.
Figure 3: The ASR results of different models on HADES using images generated at different optimization steps.
To validate the effectiveness of the optimization process for image generation discussed in Sec. 3.2, we conduct a specific experiment to examine the attack performance with the intermediately generated images with gradually increasing optimization steps. As shown in Fig. 3, the ASR results of all the comparison models consistently improve when using more optimization steps for image generation. These findings affirm the efficacy of our proposed image harmfulness optimization method in HADES.
(a) Transferability across MLLMs.
(b) Transferability across categories.
Figure 4: The evaluation results of transferability of HADES across different MLLMs (LLaVA, LLaVA-1.5 and LLaVA-1.5L) and different instruction categories (Violence, Self-Harm, Privacy, Financial, and Animal).
4.3.2 Transferability of Adversarial Attack.
To further validate the transferability of HADES across various MLLMs and harmful categories, we select Violence as the primary category for assessing cross-model transferability and LLaVA as the target model for exploring cross-category transferability. We then implement HADES utilizing $i_{\text{opt}}$ trained on a specific model/category to conduct attacks on other models/categories. The evaluation results are presented in Fig. 4. Fig. 3(a) demonstrates that HADES trained on one MLLM achieves comparable ASR on other MLLMs. Additionally, we observe that attacks between LLaVA-1.5 and LLaVA-1.5L demonstrate significant mutual transferability, likely due to their shared backbone LLMs and vision encoders. Furthermore, as illustrated in Fig. 3(b), HADES demonstrates enhanced transferability within specific harmful categories such as Self-Harm, Violence, and Animal, as well as Privacy and Financial. Such phenomena can be attributed to the similar semantic contexts shared among instructions within these categories. For instance, instructions related to Violence and Self-Harm often involve physically harmful actions such as “hitting” or “killing”, whereas those about Privacy and Financial both typically focus on abstract harmful concepts like “eavesdropping” or “forgery”.
4.3.3 Jailbreak Cases.
To better understand how our approach jailbreaks MLLMs, we analyze the successful attack cases from Gemini ProV and GPT-4V, and summarize three representative jailbreak categories, which are presented on the left side of Fig. 5. Each category is related to a distinct multimodal capability, i.e., optical character recognition (OCR), image captioning and instruction following. For the OCR category, the model explicitly recognizes the keywords from the $i_{\text{typ}}$ before following the instruction. For the Captioning category, the model describes the scenario depicted in $i_{\text{opt}}$ before following the instruction. For the Instruction Following category, the model directly follows the harmful instruction. We further calculate the proportion of these categories among all successful attack cases of Gemini ProV and GPT-4V, and illustrate the result on the right part of Fig. 5.
Figure 5: The representative cases and statistics of three harmful response types on Gemini ProV and GPT-4V. The text related to the corresponding type is underlined.
From the charts, we notice that most jailbroken cases are due to the conflicts between the general instruction following abilities of MLLMs and their harmlessness alignment, while captioning and OCR abilities also occupy considerable proportions. The results suggest that the cross-modal finetuning may impose a kind of “inverse alignment tax” on MLLMs, which improves their multimodal abilities while impairing the harmlessness alignment. Therefore, to enhance the harmlessness alignment of MLLMs, future work could consider adding more adversarial samples that consist of harmful instructions and images during the cross-modal finetuning process, which ensures MLLMs defend attacks from the image side while maintaining multimodal abilities.
5 Related Work
Harmlessness Alignment of LLMs.
Alignment refers to the process of fine-tuning LLMs with annotations based on human preferences, to ensure that the generated responses of LLMs are Helpful, Honest, and Harmless, a.k.a., the 3H principle [2]. The harmlessness alignment of LLMs has attracted extensive research attention. To evaluate the robustness of LLMs to harmful instructions, researchers employ red teaming to benchmark the safety degree of MLLMs [13, 11]. Additionally, some studies further explore the harmlessness alignment of LLMs with jailbreaking methods, either in white-box or black-box attacks. In white-box attacks, attackers have full access to the target LLMs, allowing them to calculate and backpropagate gradients from customized harmful outputs to inputs, which guides the target LLMs to generate harmful content [32, 40, 29]. In contrast, in black-box attacks, attackers have no access to the model’s parameters or gradient information, and only launch attacks through the prompting interface [34] or LLM attacker [5, 33]. Our work mainly focuses on extending the jailbreaking research from LLMs to MLLMs, which aims to enhance the robustness and alignment of MLLMs.
Multimodal LLMs.
By leveraging the excellent capacities of LLMs, the community has witnessed the emergence of several studies on MLLMs. Specifically, LLaVA [18] and MiniGPT-4 [39], which integrate an LLM, a vision encoder, and a cross-modal connector, introduce two-stage training strategies, namely cross-modal representation aligning and visual instruction tuning. Following this procedure, many open-source MLLMs have been released, by improving the multimodal performance with high-resolution and stronger LLMs [17], more pre-training corpus [3], and specially pre-training tasks [3, 6].
Harmlessness Alignment of MLLMs.
Similar to LLMs, MLLMs also suffer from malicious prompts and likely generate harmful content [24, 37]. To explore the harmfulness of MLLMs, some studies try to prob the potential harmfulness of MLLMs with benchmarks, while others employ white- and black-box attacks on MLLMs with adversarial images and prompts. However, existing benchmarks mainly focus on the effectiveness capacities of MLLMs, lacking sufficient consideration of harmlessness evaluation [10, 15, 20]. To bridge this gap, recent studies have developed a red teaming attack benchmark covering different scenarios to systematically analyze the harmlessness of MLLMs [31, 19, 16]. Similar to LLMs, existing work on jailbreaking MLLMs can be categorized into white- and black-box attacks. For white-box attacks on MLLMs, existing work [24, 26, 4, 27, 9] can be further classified into two categories according to attacking objectives: input images and visual embeddings. For input images, recent studies generate adversarial images with constraints of a harmful response set [24, 26, 9] or utilizing a teacher-forcing optimization approach [4]. For visual embeddings, Shayegani et al. [27] generates adversarial images that look harmless but are similar to the embeddings of harmful images, thereby bypassing harmful content filters. In contrast, recent work in black-box attacks jailbreak the harmlessness alignment of MLLMs by employing techniques such as system prompt attacks [35], transferring harmful information into text-oriented images [12], generating adversarial images with surrogate models [38], and maximum likelihood-based jailbreaking [21]. In our work, we first investigate whether the visual input influences the harmlessness alignment of MLLMs and then incorporate promoting engineering, white- and black-box attacks on MLLMs.
6 Conclusion
In this paper, we conducted a comprehensive empirical analysis of the harmlessness alignment of MLLMs, specifically examining the visual vulnerabilities for jailbreak. Our findings revealed that images pose significant vulnerabilities in the alignment of MLLMs: the presence of images, the cross-modal finetuning process, and the harmfulness of images all contribute to an increased propensity for MLLMs to generate harmful responses. Furthermore, we introduced HADES, a novel jailbreaking approach that hides and amplifies the harmfulness of textual instructions using meticulously crafted images. Extensive experiments have demonstrated that HADES is capable of effectively jailbreaking both open- and closed-source MLLMs. In summary, our work has presented strong evidence that the visual modality poses the alignment vulnerability of MLLMs, underscoring the urgent need for further exploration into cross-modal alignment.
For future work, we will consider improving our jailbreak strategies to more effectively jailbreak well-aligned MLLMs like GPT-4V. And we will devise cross-modal alignment methods to enhance the harmlessness alignment of MLLMs.
Appendix 0.A Defending HADES with Contrastive Harmlessness LoRA
In this section, we conduct a preliminary exploration of improving the harmlessness alignment of MLLMs. Specifically, we collect both harmful and harmless instructions related to OCR and captioning tasks, then utilize these instructions to finetune LLaVA-1.5 with LoRA. The evaluation results on HADES show that our approach can greatly reduce the harmfulness of the model’s responses, while still maintaining the model’s general multimodal capabilities.
Table 3: Evaluation results of LLaVA-1.5 and LLaVA-1.5 with contrastive harmlessness LoRA (represented by $+$ CH LoRA) on HADES and LLaVA-Bench. HADESopt and HADESadv represent the $+$Opt image and $+$Adv image setting of HADES, respectively. The better result is bolded.
Model HADESopt 0HADESadv 0LLaVA-Bench
LLaVA-1.5 79.20 89.53 63.40
$+$ CH LoRA 6.67 5.07 57.90
0.A.1 Experimental Setting
The experimental results of HADES reveal that existing MLLMs lack defenses against images containing harmful textual or visual information. The former challenge is related to the OCR abilities of MLLMs, while the latter concerns their captioning abilities. We argue that existing MLLMs struggle to resist HADES attacks because they lack corresponding alignment data in their training data for these two tasks. Therefore, we propose to mitigate such misalignment issues by collecting instructions consisting of both harmful and harmless data to finetune MLLMs.
We define each instruction as a triplet =$\mathcal{I}=\{t,i,r\}$, consisting of a textual instruction $t$, an image $i$, and an expected response $r$. For OCR tasks, we collect negative instructions $\mathcal{I_{\text{neg}}}$ from the $+$Text-to-image pointer setting in HADES, with $t_{\text{neg}}$ and $i_{\text{neg}}$ as the original instruction and image, respectively. The response $r_{\text{neg}}$ is crafted to guide the model to refuse to answer harmful instructions while maintaining OCR capability, which is designed as: “The object/concept/behavior in the image is {}. But I can not answer harmful questions.” Each $\mathcal{I_{\text{neg}}}$ is paired with a positive instruction $\mathcal{I_{\text{pos}}}$. Initially, we utilize the same image, $i_{\text{neg}}$, for $i_{\text{pos}}$. Subsequently, we transform the harmful keywords or phrases in $t_{\text{neg}}$ into a benign instruction, $t_{\text{pos}}$, using ChatGPT. We then generate a harmless response ,$r_{\text{pos}}$, also employing ChatGPT. For captioning tasks, we utilize the original visual instruction tuning data about captioning tasks of LLaVA as $\mathcal{I}_{\text{pos}}$. For $\mathcal{I}_{\text{neg}}$, we select optimized images $i_{\text{opt}}$ (mentioned in Sec. 3.2) whose harmfulness score is higher or equal to 5 as $i_{\text{neg}}$. The instruction $t_{\text{neg}}$ is set to the same as $t_{\text{pos}}$ and the response $r_{\text{neg}}$ is set to “Sorry, I can not generate harmful captions.”.
In total, we collect 2286 instructions. We shuffle all these instructions while ensuring that each negative instruction is followed by its positive counterpart. This method aims to teach MLLMs to differentiate between harmful and harmless instructions by contrasting them, thereby learning which instructions should be followed. Subsequently, we adopt these instructions to finetune LLaVA-1.5 using LoRA. The resulting LoRA is named as contrastive harmlessness LoRA.
0.A.2 Results and Analysis
To evaluate the effectiveness of our methods, we evaluate LLaVA-1.5 and LLaVA-1.5 with contrastive harmlessness LoRA on the $+$Opt image and $+$Adv image setting of HADES. Besides, we also evaluate these models on LLaVA-Bench to discuss the influence of contrastive harmlessness LoRA on the general multimodal abilities of MLLMs.
The evaluation results, detailed in Tab. 3 reveal that our contrastive harmlessness LoRA remarkably reduces the ASR of LLaVA-1.5. Specifically, its average ASR decreased from 79.20% to 6.67% in HADESopt and from 89.53% to 5.07% in HADESadv. Moreover, contrastive harmlessness LoRA doesn’t significantly impact LLaVA-1.5’s performance on LLaVA-Bench. The results suggest that finetuning MLLMs with image-related alignment data can significantly enhance their harmlessness alignment performance, while not influence other multimodal abilities.
Appendix 0.B Data Collection Pipeline
Figure 6: The harmful instruction collection pipeline.
In this section, we introduce the data collection pipeline of harmful instructions, which is presented in Fig. 6. First, we select five representative ones that are related to the visual information in the real world. Next, we adopt GPT-4 to generate 50 keywords for each of the above harmful categories, and then synthesize three harmful but distinct instructions based on the keywords. Then, we pair each instruction with a corresponding real-world image that is relevant to the harmful keywords. Specifically, we first retrieve five images from Google using the keywords as the query, and then employ CLIP to select the image that best matches the semantic representation of the keywords.
Appendix 0.C Additional Jailbreak Cases
In this section, we provide more jailbreak cases of LLaVA-1.5, GPT-4V and Gemini ProV, which are presented in Fig. 7, Fig. 8, Fig. 9 , respectively.
Figure 7: Jailbreak cases of LLaVA-1.5.
Figure 8: Jailbreak cases of GPT-4V.
Figure 9: Jailbreak cases of Gemini ProV.
Appendix 0.D ChatGPT Prompt for Harmful Instruction Generation
In this section, we present the prompt for generating harmful keywords and instructions in Fig. 10.
Figure 10: Keyword generation and instruction generation prompts.
Appendix 0.E Prompts for the Attacker and Judging Model
In this section, we present the system prompt for attacker and judging model in Fig. 11 and Fig. 12, respectively.
Figure 11: System prompt for attacker model.
Figure 12: System prompt for judging model.
Appendix 0.F Pseudo Code for Image Harmfulness Optimization
In this section, we formulate the process of optimizing image harmfulness by LLMs in Algorithm 1.
Input : Number of iterations $K$, attacker model $\mathcal{A}$, caption model $\mathcal{C}$, image generation model $\mathcal{D}$, judging model $\mathcal{J}$, system prompt template $p_{sys}$, caption prompt $p_{cap}$, initial image generation prompt $p_{0}$
Output : Optimized image set $I$
1 Initialize conversation history $h=[p_{sys}]$ Initialize optimized image set $I=\emptyset$ $i_{\text{opt}}^{0}=\mathcal{D}(p_{0})$ $I=I\cup i_{\text{opt}}^{0}$ for $k=0$ to $K-1$ do
2       $c_{k}=\mathcal{C}(i_{\text{opt}}^{k},p_{cap})$triangleright Generate an image caption
3       $[s_{k},exp_{k}]=\mathcal{J}(c_{k})$triangleright Generate the score and explanation
4       $h=h+[p_{k},c_{k},s_{k},exp_{k}]$triangleright Update the conversation history
5       $p_{k+1}=\mathcal{A}(h)$triangleright Refine the image generation prompt
6       $i_{\text{opt}}^{k+1}=\mathcal{D}(p_{k+1})$triangleright Generate a new image
7       $I=I\cup i_{\text{opt}}^{k+1}$triangleright Update the image set
8
return I
Algorithm 1 Image Harmfulness Optimization by LLMs